OpenStudy (anonymous):

PLEASE HELP!!! Last year, a poll of 1,000 voters conducted by the staff of Senator Chun found that 522 people approved the job the senator was doing. This year, a new poll of 1,000 voters showed that 423 people approve of the senator's performance. Find the percent of change in the number of voters who approve of the senators performance and identify it as an increase or decrease.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

What percentage of the people in the first poll approved?

OpenStudy (anonymous):

ummm

OpenStudy (sleepyjess):

Hint: 100/1000 = 10%

OpenStudy (anonymous):

.... 52.2%?

OpenStudy (sleepyjess):

Yes what is the percentage for the second poll?

OpenStudy (anonymous):

Wow........ 42.3?

OpenStudy (sleepyjess):

Yep so is that an increase or decrease in people that approved of the senator?

OpenStudy (mathstudent55):

You don't need the percentages fort the two polls. That is not what you are being asked. First, last year 522 approved. This year 422 approve. Is this an increase or decrease?

OpenStudy (sleepyjess):

"Find the percent of change in the number of voters"

OpenStudy (mathstudent55):

The percent of change is a comparison between two numbers. In this case, you are comparing the numbers 522 and 423. You are not comparing percentages.

OpenStudy (anonymous):

@mathstudent55 the answers are 23.4% decrease & 19% decrease and stuff. Meaning you need the percent.

OpenStudy (mathstudent55):

Yes, I understand that. The percent change is a method of comparing numbers by showing what percentage of the first number changed in becoming the second number. The answer will be a percent. I know that. That still has nothing to do with finding the percentage of people who approved the senator both times.

OpenStudy (anonymous):

Then what do I do?

OpenStudy (anonymous):

So \[99 \over 522\]

OpenStudy (anonymous):

Times 100?

OpenStudy (anonymous):

9900 over 5220... ;-;

OpenStudy (mathstudent55):

BTW, if you look at the percentages that approve the senator in the two polls, they are 52.2% and 42.3%. How will you get a 23.4% or 19% increase or decrease using these numbers?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Sorry my computer crashed. I'm back now.

OpenStudy (mathstudent55):

Above I gave you the percent change formula incorrectly. This is the correct version: \(percent ~change = \dfrac{new~number - old~number}{old ~number } \times 100 \)

OpenStudy (mathstudent55):

The correct version is ("new number" - "old number") divided by ("old number") times 100 If the result is negative, it is a percent decrease. If the result is positive, it is a percent increase.

OpenStudy (mathstudent55):

\(percent ~change = \dfrac{new~number - old~number}{old ~number } \times 100\) \(percent ~change = \dfrac{423 - 522}{522} \times 100 = \dfrac{-99}{522} \times 100 = -18.9655/% = -19\%\)

OpenStudy (mathstudent55):

The number is -18.9655..%, so you can round it to -19%. Since it is negative, it is a percent decrease.

OpenStudy (mathstudent55):

Sorry, gtg. If you have any questions, just ask. I'll try to answer next time I'm online.

OpenStudy (anonymous):

YAY I was thinking 19% decrease~ THANKS!

OpenStudy (mathstudent55):

Correct. 19% decrease is the answer.